OpenStudy (anonymous):

Use the double angle formula to solve: 10cos^2 x - 5 I have the formulas with me but I have no idea how to start this off

OpenStudy (anonymous):

is it \[10\cos^2(x) -5\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that equal to zero?

OpenStudy (anonymous):

no, that's all the problem says, that's why I'm having trouble

OpenStudy (anonymous):

Then its an expression. \[\cos(2a) = \cos^2(a)-\sin^2(a)\] so \[\cos^2(a) = \cos(2a)+\sin^2(a)\]? the issue is that you can't solve for something without an equation and an equals sign.

OpenStudy (anonymous):

okay so how do I start this problem?

OpenStudy (anonymous):

1) be very clear on what the problem asks you to do. What does it ask? maybe it is \[10\cos^2(x) = 5\]?

OpenStudy (anonymous):

It says, use the double angle formula to rewrite 10cos^2 x - 5

OpenStudy (anonymous):

Skip it and move on unless its your last problem. if it was \[10\cos^2(x) =5\] then \[\cos^2(x) = 0.5\] and \[\cos(x) = \pm 0.25\]

OpenStudy (kesumonu):

i think u can slove it like use multiple angle formula cos(2x)=2cos^2(x)-1 so in ur case 10cos^2(x)=5 => 5*(2cos^2(x))=5 => 5*(1+cos(2x))=5 => 1+cos(2x)=1 (dividing both side by 5) => cos(2x)=0 now solve for x